OSA: J Williams
Observers: D Miller
ISA: O Kuhn
Clouds and later precipitation
The night was to be split between SOUL ESM checkout during the first half and ARGOS work during the second, but clouds kept us closed.
Doug did closed dome work, testing the parameter table for SX SOUL ESM. Further tests and tweaks will require on-sky time.
Vibrations were seen on both CCD47 and LUCI1 and, with the loop open, the orientation of the vibration axis changed as LUCI1 was rotated.
Doug's log is here on the Adaptive Optics wiki at this link: https://wiki.lbto.org/AdaptiveOptics/20191208_SX_Night
- 09 Dec 2019